# UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models

### v1

@@ Abstract
Large language models require iterative updates to address challenges such as knowledge conflicts and outdated information (e.g., incorrect, private, or illegal contents). Machine unlearning provides a systematic methodology for targeted knowledge removal from trained models, enabling elimination of sensitive information influences. However, mainstream fine-tuning-based unlearning methods often fail to balance unlearning efficacy and model ability, frequently resulting in catastrophic model collapse under extensive knowledge removal. Meanwhile, in-context unlearning—which relies solely on contextual prompting without modifying the model’s intrinsic mechanisms—suffers from limited generalizability and struggles to achieve true unlearning. In this work, we introduce UniErase, a novel unlearning paradigm that employs learnable parametric suffix (unlearning token) to steer language models toward targeted forgetting behaviors. UniErase operates through two key phases: (I) an optimization stage that binds desired unlearning outputs to the model’s autoregressive probability distribution via token optimization, followed by (II) a lightweight model editing phase that activates the learned token to probabilistically induce specified forgetting objective. Serving as a new research direction for token learning to induce unlearning target, UniErase achieves state-of-the-art (SOTA) performance across batch, sequential, and precise unlearning under fictitious and real-world knowledge settings. Remarkably, in terms of TOFU benchmark, UniErase, modifying only around 3.66% of the LLM parameters, outperforms previous forgetting SOTA baseline by $\sim$ 4.01$\times$ for model ability with even better unlearning efficacy. Similarly, UniErase, maintaining more ability, also surpasses previous retaining SOTA by 35.96% for unlearning efficacy, showing dual top-tier performances in current unlearing domain. We release our code at  https://github.com/Ymm-cll/UniErase.
@@ 1 Introduction
While the large language models (LLMs) community xu2024survey ; kumar2024large ; touvron2023llama ; guo2025deepseek ; chang2024survey ; wang2025comprehensive has made significant advances in enhancing the reasoning and memorization capabilities through pretraining ma2024megalodon ; mckinzie2024mm1 ; tirumala2023d4 and post-training kumar2025llm ; gan2024towards ; huang2024billm ; tie2025survey techniques, the complementary challenges of enabling precise and efficient unlearning remain crucial to address privacy, ethical and security concerns lu2024eraser ; zhang2024safe ; xu2024machine ; si2023knowledge . The heterogeneous nature of training data sources and massive parameters of LLMs introduce significant difficulties in controlling the generation of sensitive, private, or illegal contents yao2024survey ; shi2024large . These privacy-related risks and potential misuse cases have made all individual entities very cautious when deploying LLMs – let alone large enterprises and governmental organizations he2024emerged ; wang2025comprehensive ; yu2025survey .
Striking an optimal balance between model ability and unlearning efficacy remains fundamentally challenging geng2025comprehensive ; qu2024frontier . The core objective of unlearning is to enable LLMs trained on trillion-token corpora to completely forget specific data subsets while preserving retaining data and their general capabilities si2023knowledge ; maini2024tofu . Instead of prohibitively expensive retraining, fine-tuning (FT) via carefully designed loss functions is the current mainstream methods, which can be broadly categorized into two types: (I) Targeted unlearning involves deliberately erasing certain data or learned associations in LLMs to modify their outputs in a controlled and specified manner zhang2024negative ; wei2021finetuned ; rafailov2023direct ; maini2024tofu ; yuan2024closer . (II) Untargeted unlearning shifts the responses away from the original outputs but without specifying a particular direction maini2024tofu ; zhang2024negative ; yuan2024closer ; wang2025uipe ; wang2025rethinking . Typically, this line employs techniques such as gradient ascent maini2024tofu and negative preference optimization zhang2024negative to steer the model away from knowledge in the forget set. Notably, in-context unlearning (ICU) is a complementary paradigm that approximates unlearning through prompt engineering alone, without modifying the model’s parameters pawelczyk2023context ; takashiro2024answer ; bhaila2024soft .
@@
Figure 1: Our UniErase achieves comprehensive and balanced unlearning (Left) while maintains consistently high LLM capability (Middle), delivering rapid processing and high precision (Right).
However, the observed empirical evidence gaps persist across the above methods, indicating non-trivial challenges in current unlearning frameworks geng2025comprehensive ; maini2024tofu ; yuan2024closer . Specifically, ICU-based methods suffer from poor generalizability, robustness, and adaptability pawelczyk2023context —for example, they can only handle unlearning in classification tasks and fail to achieve true forgetting. Meanwhile, FT-based approaches struggle to strike a balance between unlearning efficacy and model ability maini2024tofu ; qu2024frontier , with other problems like over-unlearning and unpredictable behaviors yuan2024closer . Empirical analysis reveals a critical limitation, for batch unlearning, the existing best method in forgetting loses 80.28% of the LLM’s general ability, while the retaining best only forgets around half of the required data.
Generally, when predicting the next token’s probability distribution, LLMs are heavily influenced by the preceding token sequences. Specifically, in the safety domain, the affirmative or negative tone of a model’s initial response can significantly alter the success rate of replying to harmful queries zou2023universal ; qi2024safety ; dong2024attacks . Then, for the unlearning scenario, a natural but non-simple question arises: Can an unlearning token, denoted as [UNL], redirect the LLM’s “cognition” of the forgetting set $\mathcal{D}_{f}$ toward an unlearning space while preserving the response direction for the retaining set $\mathcal{D}_{r}$?
To this end, we propose UniErase, a universal unlearning method that balances unlearning efficacy and model ability. Technically, UniErase mainly consists of two parts: unlearning token and unlearning editing. Inspired by meta-tokens li2021prefix ; lester2021power , UniErase first learns an unlearning token (denoted as [UNL]) on the forgetting set $\mathcal{D}_{f}$ and ensures other LLM parameters remaining unaffected by freezing them. To learn [UNL], we only introduce two new vectors in the embedding and unembedding matrices (projecting hidden states into next-token probability distribution). Through optimization on these two vectors, the LLM is guided to output “I don’t know”-like responses when [UNL] following input knowledge (in $\mathcal{D}_{f}$) queries. Notably, without [UNL], the LLM still responds $\mathcal{D}_{f}$ and the retaining set $\mathcal{D}_{r}$ normally. Thus, to build the connection between [UNL] and $\mathcal{D}_{f}$, we further propose unlearning editing, which modifies a minimal amount of model parameters to steer $\mathcal{D}_{f}$ into the “unlearning space” via [UNL] and preserves knowledge in $\mathcal{D}_{r}$. Specifically, by adjusting a few parameters in specific Transformer layers, for the knowledge queries in $\mathcal{D}_{f}$, we alter the first token of their responses to [UNL], leveraging its properties to ensure the model forgets these knowledge. In a parallel vein, we maintain the knowledge associations in $\mathcal{D}_{r}$ to keep the LLM’s general ability.
@@
We conduct extensive experiments to validate the effectiveness of our UniErase. In fact, we consider fictitious and real-world knowledge in batch, sequential and precise unlearning scenarios for the Llama series of LLMs. Evaluating via multi-dimensional metrics on TOFU benchmark, UniErase significantly outperforms the previous SOTA forgetting baseline, achieving a 4.01$\sim$ 4.01$\times$ 4.01 × improvement in model ability while demonstrating better unlearning. Additionally, compared with prior retaining SOTA, UniErase maintains superior model ability and is 35.96% higher in forgetting efficacy.
@@
  • Brand-new Paradigms. Building upon UniErase, we propose a novel paradigm that achieves efficient unlearning by directly modifying model parameters instead of multi-round fine-tuning, significantly expanding the scope of future research in the current unlearning technology stack.
  • Balanced Trade-off. Compared to previous SOTA methods, UniErase achieves more thorough unlearning and better general ability preserving, boosting the practical usability of LLM unlearning.
  • Generalized Scenarios. UniErase achieves the best performances across batch, sequential and precise unlearning for fictitious and real-world knowledge, covering nearly all unlearning cases.
@@ 2 Related Works
Machine Unlearning. The concept of machine unlearning from traditional models bourtoule2021machine ; chen2022graph ; nguyen2022survey is emerging as a rising approach for LLM alignment research liu2025rethinking ; thaker2024position . Its primary goal is to enable LLMs to forget a subset $\mathcal{D}_{f}$ of training data $\mathcal{D}$ (e.g., privacy or harmful knowledge) and maintain the abilities on a retain set $\mathcal{D}_{r}\subset\mathcal{D}$, without the high cost of retraining geng2025comprehensive . Current mainstream approaches relying on fine-tuning and designing various loss functions for different objectives (forget $\mathcal{D}_{f}$ and retaining $\mathcal{D}_{r}$). Specifically, GD liu2022continual reduces the probability of generating responses to $\mathcal{D}_{f}$ by ascending gradients, and introduces another loss to constrain the deviation between the unlearned model and the original one liu2022continual . Meanwhile, NPO zhang2024negative , based on preference optimization rafailov2023direct , achieves unlearning by solely using $\mathcal{D}_{f}$ as negative preference, ignoring the positive term. Other works, such as RMU huu2024effects and LUNAR shen2025lunar , employ steering-vector-like approaches cao2024personalized to forcibly modify hidden states (but still via fine-tuning) and redirect $\mathcal{D}_{f}$ toward inability space. Additionally, SPULbhaila2024soft makes preliminary attempts in unlearning by adding soft prompts li2021prefix ; lester2021power during inference to manipulate model responses, but without modifying parameters to achieve “essential forgetting”.
Model Editing. LLMs may contain outdated or even incorrect information huang2025survey ; tonmoy2024comprehensive . However, like unlearning, retraining to remove is costly, while fine-tuning overfits with limited updates. Thus, the light and precise model editing wang2024knowledge ; he2024llms is proposed for truthfulness huang2024can , and safety li2024precision ; chen2024can ; li2024badedit . Specifically, interpretability study treats LLMs’ two-layer Multilayer Perceptrons (MLP) as lookup tables for storing knowledge geva2020transformer ; zhao2021non . So Early methods like ROME meng2022locating and MEMIT meng2022mass introduce the locate-then-edit paradigm by solving closed-form optimization on the MLP’s down-projection matrix. AlphaEdit fang2024alphaedit further efficiently preserves other knowledge by projecting updates into the null space of the original knowledge representation. AnyEdit jiang2025anyedit decomposes edits into sub-operations for more complex knowledge. However, editing itself cannot accomplish unlearning directly: it requires knowledge represented in (subject, relation, object) triples and updates the object to a new value, but there exists no object token that corresponds to the abstract concept of unlearning. Our UniErase introduces unlearning-specific editing that integrates with existing methods.
@@ 3 Preliminaries
Notations. We consider an LLM as a probabilistic language model with parameters $\theta$, denoted as $\pi_{\theta}$. The target set for forgetting and the set of retained knowledge are represented as $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, where typical elements of both are question and answer pairs in the form of $d=(q,a)$. In addition, we denote the set of real numbers as $\mathbb{R}$, and the set of real number tensors with dimensions $(d_{1},...,d_{n})$ as $\mathbb{R}^{d_{1}\times...\times d_{n}}$. The symbol $\oplus$ represents the concatenation operation of token sequences.
Unlearning Target. For an LLM $\pi_{\theta}$ trained with dataset $\mathcal{D}$, unlearning process aims to make the model forget the contents in $\mathcal{D}_{f}\subset\mathcal{D}$, such that it behaves as if it were trained solely on $\mathcal{D}\setminus\mathcal{D}_{f}$. In a parallel vein, unlearning must preserve the model’s knowledge on other datasets and its general capabilities. Similar to the trade-off between safety and helpfulness in LLM optimization varshney2023art , unlearning involves a balance between unlearning efficacy and model ability, formulated as:
@@
where “$\mathrm{Forget}$” and “$\mathrm{Ability}$” are the standards or metrics for unlearning efficacy and model ability.
Mainstream Unlearning Strategies. To achieve the goal in Eq 1, mainstream unlearning methods design distinct forgetting/retaining losses $l_{f}/l_{r}$ on $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, sometimes using the original model $\pi_{\theta}^{\text{ref}}$ as a reference. We unify their loss designs as follows, with $\beta$ and $\gamma$ as the weights:
@@
In Eq 2, the forgetting term is designed to make the model forget the content on $\mathcal{D}_{f}$, while the retaining term aims to preserve the knowledge on $\mathcal{D}_{r}$. However, current methods typically select the neighboring knowledge of $\mathcal{D}_{f}$ as $\mathcal{D}_{r}$, and can not encompass diverse general knowledge due to high time cost. In Appendix B, we introduce the specific forms of various $l_{f}$ and $l_{r}$ in detail.
4 Methodology
In this section, we initially employ mathematical logic to expound upon the fundamental principles of UniErase ($\triangleright$ Sec 4.1), as demonstrated in Figure 2. Then, we elaborate on the methodology of training an unlearning token via incorporating a minimal number of parameters ($\triangleright$ Sec 4.2). Additionally, we discuss the process of precisely modifying parameters for the forgetting targets ($\triangleright$ Sec 4.3), aiming at achieving an optimal equilibrium between high unlearning efficacy and low ability loss ($\triangleright$ Sec 4.4).
@@ 4.1 Unlearning Logical Chain
Given an LLM $\pi_{\theta}$, for an input token sequence $q=[x_{1}x_{2}...x_{n}]$, we assume the output token sequence is $a=[y_{1}y_{2}...y_{m}]$. Then we abstract this generation process as mathematical logic derivations:
@@
Proposition 1. The Unlearning Token is a novel token trained on $\mathcal{D}_{f}$, designed to constrain $\pi_{\theta}$ such that its subsequent token generation redirects targeted concept forgetting or induced unawareness. Specifically, for $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$, the unlearning token [UNL] is defined as follows:
@@
$x_{1}x_{2}...x_{n}\oplus\mathrm{[UNL]}\xRightarrow{\pi_{\theta}}y_{\text{idk}},\quad y_{\text{idk}}\in\mathcal{D}_{\text{idk}},$ (4)
where $\mathcal{D}_{\text{idk}}$ is the set of token sequences that express forgetting or ignorance (e.g., “I do not know”).
@@
Building upon Proposition 1, we have learned the [UNL] token. However, when only $q$ is provided as input, the model still generates original normal response rather than $y_{\text{idk}}$. Therefore, we need to modify model parameters so that for any $q\in\mathcal{D}_{f}$, the first token of the generated response is [UNL], thereby internalizing “forgetting $q$ with [UNL]” as part of the model’s knowledge or “cognition”. Model editing, as an evolving approach to edit model knowledge without ability degradation fang2024alphaedit ; meng2022locating , basically meets our unlearning requirements if adapted properly. To this end, we propose:
@@
Proposition 2. Unlearning Editing utilizes model editing to modify only a small set of parameters ($|\Delta\theta|\ll|\theta|$), enabling the model to forget specified intrinsic knowledge. For $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$ with the obtained [UNL], the parameter $\theta$ is updated with $\Delta\theta$, formally defined as:
$\theta\leftarrow\theta+\Delta\theta,\quad|\Delta\theta|\ll|\theta|,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}.$ (5)
Derivation: Eq 4 ensures that [UNL] can direct the input to the unlearning space (generating responses in $\mathcal{D}_{\text{idk}}$) when used as a suffix, while Eq 5 changes the next token of the normal response of the input sequence to [UNL]. Based on these two propositions, we can derive the following Unlearning Logical Chain to adjust model parameters for achieving our UniErase unlearning:

### v2

# UniErase: Towards Balanced and Precise Unlearning in Language Models

@@ Abstract
[…]
@@ 1 Introduction
While the Large Language Models (LLMs) community (Guo et al., 2025; Chang et al., 2024; Wang et al., 2025a) has made significant advances in “learning” general abilities and domain-specific knowledge via pretraining and post-training (Kumar et al., 2025; Tie et al., 2025). Meanwhile, an equally crucial research direction is the complementary concept of LLM “unlearning” (Liu et al., 2025; Geng et al., 2025), which serves to address critical issues related to hallucination (Huang et al., 2025), privacy (Yan et al., 2024), and safety (Wang et al., 2025a)—including updating outdated knowledge, removing private information, and eliminating harmful contents (Lu et al., 2024; Zhang et al., 2024c; Xu, 2024). The core objectives of ideal unlearning is to enable LLMs, trained on trillion-token corpora, to only forget a specific data subset (the forgetting set) without compromising their general knowledge (the retaining set) and capabilities (Si et al., 2023; Maini et al., 2024).
Given the prohibitive computational cost of retraining LLMs from scratch while excluding the forgetting set, fine-tuning (FT) techniques has emerged as the predominant unlearning implementation (Maini et al., 2024; Yuan et al., 2024). Concretely, FT-based unlearning can be broadly categorized into two paradigms: (I) Targeted unlearning deliberately modifies LLMs’ outputs of the forgetting set in controlled and specified manners, such as “I don’t know”-like expressions (Wei et al., 2021; Rafailov et al., 2023); (II) Untargeted unlearning shifts the responses away from the original outputs but without specifying a particular direction, like irrelevant answers (Maini et al., 2024; Zhang et al., 2024b). These two paradigms both employ carefully designed loss functions with distinct objectives for the forgetting set (forgetting loss) and retaining set (retaining loss), respectively, thereby enabling knowledge erasure and retention (Yuan et al., 2024; Wang et al., 2025c; b).
However, fine-tuning inherently requires sufficient data volume to achieve effective optimization without overfitting, and the forgetting loss and retaining loss present competing objectives (Yuan et al., 2024; Geng et al., 2025). Besides, current FT-based unlearning, due to cost considerations, limits the retaining set to distributions near the forgetting set, which cannot represent the broad LLM knowledge (Maini et al., 2024). Consequently, these render two critical goals fundamentally challenging: ❶ precise unlearning for a small or even single-entry forgetting set, or ❷ balanced unlearning that concurrently preserves general abilities and knowledge while ensuring high unlearning efficacy on the forgetting set (Veldanda et al., 2024; Qu et al., 2024). Our empirical experiments across 8 baselines validate the second dilemma: for batch unlearning, the best-forgetting method loses 80.28% of the general abilities, while the best-retaining baseline only forgets $\sim$ half of the target data.
@@
Figure 1: Our UniErase achieves the most balanced unlearning performances (Left) and maintains consistently high general capabilities (Middle), delivering rapid processing and high precision (Right).
In this paper, we aim to tackle these two critical issues for LLM knowledge unlearning, especially the balance challenge. To this end, we propose UniErase, a novel unlearning paradigm that balances unlearning efficacy and model abilities with dual-high performances, while supporting effective and efficient precise unlearning. Technically, UniErase consists of two innovative techniques: the Unlearning Token and the Unlearning Edit (Udit). We first introduce the unlearning token that concretizes the concept of forgetting into a tangible entity and points to a representational space that encodes unlearning semantics. Specifically, the unlearning token directs the autoregressive prediction process to generate predefined forgetting responses for any input sequence that terminates with this token. To obtain it without affecting any other generation, we create and optimize a new meta token (Li & Liang, 2021; Lester et al., 2021) only in the embedding space of the LLM, with other parameters frozen. Building upon this, we further propose Udit, a data-volume-independent method (therefore supporting precise unlearning) that directly modifies model parameters to establish associations between the forgetting set and unlearning token, thus realizing unlearning via its directing property. More importantly, Udit employs the null space projection technique (Fang et al., 2024) to ensure parameter updates remaining orthogonal to the LLMs’ existing knowledge representations, effectively preserving the retaining set and even general capabilities.
@@
In contrast to FT-based unlearning, UniErase pioneers the modeling of LLM unlearning as a knowledge editing problem. We solve the problem that current LLM editing frameworks only support entity concept editing (Wang et al., 2024; Zhang et al., 2024a) via the unlearning token, and further propose Udit to truly achieve precise and balanced LLM unlearning. To validate the effectiveness of UniErase, following previous works (Yuan et al., 2024; Zhang et al., 2024b), we conduct extensive experiments on different scales of the Llama-3 (Dubey et al., 2024) LLMs. Actually, with 8 baselines, we consider both fictitious and real-world knowledge in batch, sequential and precise unlearning scenarios (as illustrated in Figure 1). Evaluating via multi-dimensional metrics on the TOFU (Maini et al., 2024) benchmark, UniErase significantly outperforms the previous best-forgetting baseline, attaining $4.01\times$ performances in maintaining general knowledge and abilities while demonstrating better unlearning. Additionally, compared with the best-retaining baseline, UniErase preserves superior LLM abilities and is 35.96% higher in unlearning efficacy.
@@
  • Brand-new Paradigm. Our proposed UniErase represents a novel unlearning paradigm that exhibits outstanding performances by directly modifying LLM parameters instead of multi-round fine-tuning, significantly expanding the scopes of future research in the unlearning community.
  • Dual-high Balance. UniErase achieves more thorough unlearning with better retention for general knowledge and abilities, boosting the practical usability of LLM unlearning.
  • Generalized Scenarios. UniErase performs superbly across batch, sequential and especially precise unlearning for fictitious and real-world knowledge, covering diverse unlearning tasks.
@@ 2 Related Works
Machine Unlearning. The concept of machine unlearning (Bourtoule et al., 2021) from traditional models (Chen et al., 2022; Nguyen et al., 2022) is emerging as a rising research topic for LLMs (Liu et al., 2025; Thaker et al., 2024). Its primary goal is to enable LLMs to forget a subset $\mathcal{D}_{f}$ (e.g., privacy or harmful knowledge) of the training data $\mathcal{D}$ and maintain the knowledge on a retaining set $\mathcal{D}_{r}\subset\mathcal{D}$, without the high cost of retraining (Geng et al., 2025). Mainstream approaches relying on the fine-tuning techniques and designing various loss functions for different objectives to simultaneously forget $\mathcal{D}_{f}$ and retain $\mathcal{D}_{r}$. For example, GD (Liu et al., 2022) reduces the probability of generating outputs in $\mathcal{D}_{f}$ by ascending gradients, and introduces another loss to constrain the deviation. Meanwhile, NPO (Zhang et al., 2024b), inspired by preference optimization (Rafailov et al., 2023), realizes unlearning by solely using $\mathcal{D}_{f}$ as negative preferences, ignoring the positive terms. Other works, such as RMU (Huu-Tien et al., 2024) and LUNAR (Shen et al., 2025), employ steering-vector-like approaches (Cao et al., 2024) to forcibly modify hidden states and redirect $\mathcal{D}_{f}$ toward the inability space. Additionally, SPUL (Bhaila et al., 2024) makes preliminary attempts in unlearning by adding soft prompts (Li & Liang, 2021; Lester et al., 2021) during inference to manipulate model responses, but without modifying parameters to achieve essential forgetting.
Model Edit. LLMs may contain outdated, incorrect or even harmful information (Huang et al., 2025; Tonmoy et al., 2024). However, similar to unlearning, retraining for knowledge updates is costly, while fine-tuning overfits for precise scenarios. Thus, the model edit techniques (Wang et al., 2024; He et al., 2024) are proposed for truthfulness (Huang et al., 2024), and safety (Chen et al., 2024; Li et al., 2024). Early methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) introduce the locate-then-edit paradigm by modifying the down-projection matrices in the LLMs’ Multi-layer Perceptron (MLP) module. AlphaEdit (Fang et al., 2024) further preserves other knowledge via the null space projection operation. However, recent unlearning surveys like  (Liu et al., 2025) have highlighted challenges including undefined edit objectives if directly applying editing for unlearning. In fact, editing itself targets at knowledge represented in the (subject, relation, object) triple formats and modifies the object to a new value (Meng et al., 2022a; Wang et al., 2024; Li et al., 2025), yet no single object token exists for the abstract concept of unlearning. Our UniErase addresses these fundamental issues by introducing Udit with the unlearning token.
@@ 3 Preliminaries
Notations. We refer to an LLM with parameters $\theta$ as $\pi_{\theta}$. The target knowledge for the forgetting set and retaining set are represented as $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, where typical elements of both are question $q$ and answer $a$ pairs in the form of $d=(q,a)$. In addition, we denote the set of real numbers as $\mathbb{R}$, and the set of real number tensors with dimensions $(d_{1},...,d_{n})$ as $\mathbb{R}^{d_{1}\times...\times d_{n}}$.
Unlearning Target. For an LLM $\pi_{\theta}$ trained with dataset $\mathcal{D}$, unlearning aims to make the model forget the contents in $\mathcal{D}_{f}$ as if it were trained solely on $\mathcal{D}\setminus\mathcal{D}_{f}$. In a parallel vein, unlearning must preserve the model’s knowledge in $\mathcal{D}_{r}$ and even broader knowledge with general capabilities. Similar to the trade-off between harmless and helpfulness in LLM safety alignment (Varshney et al., 2023), unlearning involves a balance between the unlearning efficacy and model ability, formulated as:
@@
where $\mathrm{Forget}$” and $\mathrm{Ability}$” are the standards or metrics for unlearning efficacy and model ability.
Mainstream Unlearning Paradigms. To achieve the goal in Eq. 1, current FT-based unlearning design diverse forgetting losses $l_{f}$ and retaining losses $l_{r}$, respectively, sometimes using the original model $\pi_{\theta}^{\text{ref}}$ as a reference. We unify their loss designs as follows, with $\beta$ and $\gamma$ as trade-off weights:
@@
In Eq. 2, the forgetting term is designed to make the model forget the contents on $\mathcal{D}_{f}$, while the retaining term aims to preserve the knowledge on $\mathcal{D}_{r}$. Current methods typically select $\mathcal{D}_{r}$ to be the neighboring knowledge of $\mathcal{D}_{f}$, which can not encompass diverse general knowledge and abilities. In Appendix B, we introduce the specific forms of various $l_{f}$ and $l_{r}$ in detail.
4 Perform Unlearning Edit with Unleanring Token
In this section, we first introduce the Unlearning Logical Chain to expound upon the fundamental principles of UniErase ($\triangleright$ Section 4.1), as demonstrated in Figure 2. Then, we propose the unlearning token and elaborate on the techniques to obtain it via incorporating a minimal number of parameters ($\triangleright$ Section 4.2). Subsequently, we introduce Udit to modify parameters for the unlearning targets ($\triangleright$ Section 4.3), achieving balanced unlearning performances while supporting precise unlearning.
@@ 4.1 Unlearning Logical Chain
Given an LLM $\pi_{\theta}$, for an input token sequence $q=[x_{1}x_{2}...x_{n}]$, we assume the output token sequence is $a=[y_{1}y_{2}...y_{m}]$. Then we abstract this generation process as a mathematical logic derivation:
@@
Proposition 1. The Unlearning Token (denoted as [UNL]) is a novel token, designed to direct the LLM’s subsequent token generation to specific forgetting expressions. We refer to the token concatenation operator as $\oplus$. Then, for any $(q,a)\in\mathcal{D}$, [UNL] satisfies the following property:
@@
$x_{1}x_{2}...x_{n}\oplus\mathrm{[UNL]}\xRightarrow{\pi_{\theta}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}\quad\land\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{\theta}}a,$ (4)
where operation $a\land b$ means that both $a$ and $b$ should be satisfied and $\mathcal{D}_{\text{idk}}$ contains different token sequences that express the semantics of forgetting or ignorance. Specifically, Eq. 4 stipulates that the newly acquired unlearning token should exclusively direct the model toward the forgetting semantic space when employed as a suffix, while preserving normal knowledge retrieval capabilities otherwise.
@@
In Proposition 1, we have defined the [UNL] meta token. However, when only $q$ is provided as input, the LLM still generates original normal responses rather than $y_{\text{idk}}$. To realize unlearning, we need to modify model parameters so that: for any $q$, its next token prediction is [UNL], thereby internalizing “forgetting $q$ with [UNL]” as the LLM’s inherent knowledge. To this end, we propose:
@@
Proposition 2. Unlearning Editing (Udit) modifies only a small set of parameters $\Delta\theta$, enabling the LLM to forget specified knowledge. For $\forall(q,a)\in\mathcal{D}_{f}$ and $\forall(q^{\prime},a^{\prime})\in\mathcal{D}\setminus\mathcal{D}_{f}$, Udit ensures that: